Could you help me write the slope-intercept equation for the line that passes through (-1, -4) and (4, 11)? Thank you.
Answer by mananth(16946)

x1 y1 x2 y2
-1 -4 4 11

slope m =(y2-y1)/(x2-x1)
(11+4)/( 4+1)
(15/5)
m=3

Plug value of the slope and point(-1,-4 ) in
Y = m x + b
-4.00=-3 +b
b=-4-3
b=-1
So the equation will be
Y =3x-1
m.ananth@hotmail.ca